OpenStudy (anonymous):

What is the solution set for |–2x + 8| = 12

OpenStudy (anonymous):

CaseI -2x + 8 = 12 -2x = 4 x = -2 Case II - (-2x+8) = 12 -2x + 8 = -12 -2x = -20 x=10

OpenStudy (anonymous):

\[- 2x + 8 = \pm 12\] Now, we take \[-2x + 8 = + 12\] We get x = -4 When we take \[- 2x + 8 = -12\] We get x = 10. Hence the answer is x = -4 or x= 10.